Problem 6 Calculate the test statistic. (A z-score and a t-score are examples of test statistics.) In order to present your final results with a numerical accuracy of 0.01, it will be necessary to perform your intermediate calculations with an accuracy of O.001 Part 1 (696) The cost of a daily newspaper varies from city to city. A study was done to test the claim that the average cost of a daily newspaper is $0.50 Ten cities were surveyed; the average cost was $0.50 with a standard deviation of $0.06 Part 2 (696) Your statistikcs instructor claims that 60 percent of the students who take an Elementary Statistics class go through life feeling more enriched. For some mysterious reason, most people Elementary Statistics students and find that 35 feel more enriched as a result of the class. dont believe this. You decide to check this out on your own. You randomly survey 50 former
Answer #1

part 1)

t = (xbar-µ)/(s/√n) = (0.50-0.50)/(0.06/√10) = 0.00

part 2)

Ho :   p =    0.6
H1 :   p ╪   0.6
      
Level of Significance,   α =    0.05
Number of Items of Interest,   x =   35
Sample Size,   n =    50
      
Sample Proportion ,    p̂ = x/n =    0.7
      
Standard Error ,    SE = √( p(1-p)/n ) =    0.06928
      
Z Test Statistic =    Z = ( p̂-p)/SE =    1.4434

p-Value   =   0.148914673      
Conclusion:     p value>α  don't reject null hypothesis

hence, there is no enough evidence to reject the claim at 0.05 significance level